• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Marrone

    Log entry time 23:54:09 on June11,2005

    Entry number 145903

    keyword=Shift Summary

    16:00 Shift started with MCC tuning for better energy spread with OTR in.

    17:10 Energy spread 3.9 e-5, OTR out.
    Started taking data
    17:10-17:42 Run 4524, MCC has injector problem, run ended prematurely

    18.00 Generally we have beam trips. The beam has also a bad energy resolution.

    19.25 Run 4526 We changed the Prescale factor T5 from 70 to 30 (see Halog entry)

    22.00 Beam is off. They are trying to identify the problems.
    24.00 The situation is not improved. They were lookig for the experts
    on call. MCC will give update as soon as possible.


    List of Runs:
    #Run, #Events, Comment
    4524, 1 M, Good
    4525, 300 k, Bad energy resolution of the beam
    4526, 1 M, Good
    4527, Coda crashed
    4528, 1 M, Good
    4529, 1 M, Good
    4530, 1 M, Good
    4531, 1 M, Good
    4532, 400 k, Good -> beam is off